# Large Language Models Powered Context-aware Motion Prediction

### v1

@@ Abstract
Motion prediction is among the most fundamental tasks in autonomous driving. Traditional methods of motion forecasting primarily encode vector information of maps and historical trajectory data of traffic participants, lacking a comprehensive understanding of overall traffic semantics, which in turn affects the performance of prediction tasks. In this paper, we utilized Large Language Models (LLMs) to enhance the global traffic context understanding for motion prediction tasks. We first conducted systematic prompt engineering, visualizing complex traffic environments and historical trajectory information of traffic participants into image prompts—Transportation Context Map (TC-Map), accompanied by corresponding text prompts. Through this approach, we obtained rich traffic context information from the LLM. By integrating this information into the motion prediction model, we demonstrate that such context can enhance the accuracy of motion predictions. Furthermore, considering the cost associated with LLMs, we propose a cost-effective deployment strategy: enhancing the accuracy of motion prediction tasks at scale with 0.7% LLM-augmented datasets. Our research offers valuable insights into enhancing the understanding of traffic scenes of LLMs and the motion prediction performance of autonomous driving.
@@ I INTRODUCTION
Motion prediction is one of the most important tasks in the field of autonomous driving [1, 2, 3], which predicts the motion statuses of nearby agents by jointly considering nearby agents and road maps. This information will assist the decision module in making a more robust and safer driving decision. Therefore, this field already features a plethora of datasets and public competitions. For instance, since 2021, Waymo has been organizing competitions in motion prediction [^1] https://waymo.com/open/challenges/2021/motion-prediction/, attracting models and algorithms that have won championships, such as MTR [1], MTR++ [2], and MGTR [4]. There are three classes of motion prediction methods: goal-based methods [5, 6], direct-regression methods [7, 8], and methods that take the best of both[1, 2]. Although methods that take the best of both can balance results and computing resources greatly, they still encode information about agents and maps first and then decode to obtain final results. The methods above all lack a comprehensive understanding of overall traffic semantics.
@@
Large Language Model (LLM) is popular in the field of autonomous driving, as not only the Language Model but also the World Model with common sense [9, 10, 11]. After the release of OpenAI’s GPT series of large language models, many works that use GPT to assist driving have been emerging [12, 13, 14, 15, 16]. These efforts leverage the inferential capabilities of large language models or through experimentation, aiding autonomous driving algorithms in better perception [17] and planning [12, 14, 15], and even facilitating end-to-end decision-making [18, 13]. However, teaching LLM to understand the interactive context among various transportation participants in diverse transportation scenarios is still a challenging task. Although many have applied vision-language models to traffic scenarios, their performance heavily relies on datasets annotated from a first-person perspective [19, 20, 21]. For abstract, information-overwhelming, and bird’s-eye view traffic scenarios, there is a lack of relevant datasets and understanding methodologies.
Figure 1: Context-aware Motion Prediction Based on LLMs. We first visualize structured information in the motion prediction dataset, GPT4-V then understand the scenario we provided via visualized image and well-designed prompt and output transportation context information. This information will be used to augment traditional motion prediction algorithms.
@@
In this paper, we propose a new method that enhances the traffic context understanding of motion prediction models using LLMs to make more accurate predictions. To enable an unfine-tuned LLM—GPT4 with vision (GPT4-V) to understand traffic context and output the necessary context information for motion prediction, we visualize the vector map data and historical data of traffic participants as a Transportation Context Map (TC-Map) to serve as an image prompt and propose a corresponding text prompt design. After successfully obtaining traffic context information generated by the LLM, we integrated this information into a classical motion prediction algorithm—MTR [1]. The results indicate that this context information effectively improves the accuracy of motion prediction. Additionally, considering the cost of LLMs, we also propose a cost-effective deployment strategy: by utilizing 0.7% LLM-enhanced datasets, we can empower motion prediction performance at scale.
@@
  • We systematically designed and conducted prompt engineering to enable an unfine-tuned GPT4-V to comprehend complex traffic scenarios involving multiple participants and to output context information such as intention, affordance, and scenario;
@@ II Method
Figure 2: Details of Transportation Context Generation Prompt.
@@
II-A Get Transportation Context from GPT4-V
Translating images directly into motion poses is significantly challenging [22]. We obtain transportation context through methods based on LLMs with vision, defining the interested agent in the TC-Map as the ego-agent, which includes three categories: vehicle, pedestrian and cyclist. As shown in Fig. 2, the TC-Map designed based on the WOMD serves as the Rasterized Image prompt, and together with the Text prompt forms the Transportation Context Generation Prompt (TCGP). Through GPT4-V, this input yields outputs such as Situation Understanding, Reasoning, Intentions of ego-agent, Affordances [23], and Scenario Types. Additionally, during the prompt design process, we summarized 6 prompt design suggestions as shown in Fig. 3.
@@ II-A1 TC-Map design suggestions
Corresponding to Prompt Design Suggestion 1 and 2 in Fig. 3. Suggestion 1 aims to reduce confusion in the volume of information. Inspired by the MTR Map Collection module, we introduced TC-Map Crop. Additionally, during the initial testing of GPT4-V, it was found that embedding a large amount of data information in the Text prompt, such as 11 frames of historical trajectory in JSON format, easily confuses GPT4-V and increases the uncertainty of intent. So the extent of TC-Map is cropped based on the type of agent (vehicle: pedestrian: cyclist = 6:4:3), limiting the ego-agent to focus only on the appropriate surrounding area scenes, reducing the impact of irrelevant factors on the prediction of Intention. Suggestion 2 aims to minimize confusion in direction identification. The diversity of scenarios leads to a variety of headings for the ego-agent, a common map guide icon is selected to normalize the heading of the ego-agent to “North”.
@@ II-A2 Text Prompt design suggestions
Corresponding to Prompt Design Suggestion 3 and 4 in Fig. 3. Suggestion 3 proposes Examples, divided into Positive and Negative categories. Positive Examples adhere to the randomness and diversity of intent generation. Negative Examples are bad cases selected from GPT4-V’s responses during the quantitative experiment of prompts, focusing on scenarios with high misjudgment rates and uncommon, easily overlooked Corner-Cases. Additional Error Analysis is provided and placed closer to the end of the Text prompt. Suggestion 4 consists of two sentences. The first sentence discusses using rules to determine the lane type of the ego-agent’s driving lane in TC-Map, the second sentence provides explanations for the added lane types, they assist GPT4-V in excluding impossible intents under the premise that the ego-agent follows traffic rules, based on traffic driving regulations.
@@ II-A3 Detailed Information Supplement
Corresponding to Prompt Design Suggestion 5 in Fig. 3. To fully leverage the scene information from WOMP dataset, such as road_edges and crosswalks, we optimize the representation of different types of scenario information in TC-Map by adding more indicative head arrows and visual labels for agents. During this phase, we also incorporate the speed of the ego-agent, as well as the speed, relative positions, and distances of the surrounding three agents.
@@ II-A4 Multi-task Outputs Requirement
Corresponding to Prompt Design Suggestion 6 in Fig. 3, we designed seven types of Intentions for the ego-agent over the next 8 seconds. STRAIGHT-LEFT and STRAIGHT-RIGHT are considered scenarios where future ego-agent merging might occur, indicating complex intentions. Predicting these intentions often requires GPT4-V to utilize rasterized images, the ego-agent’s own position, heading, and speed, the state of surrounding vehicles, and the direction of nearby lane lines predicting these two types of intentions highly challenging. Additionally, based on an understanding of complex scenes, we designed five types of Affordances and four types of Scenario types, collectively providing a rich, human-like transportation context information to enhance the accuracy of the motion prediction task.
@@ II-B1 Motion Transformer
We choose Motion TRansformer (MTR), SOTA of WOMD Challenge 2022, as our base model. MTR is developed based on the architecture of Transformer [24]. It first encodes the scenario information including agents and map roads, and then obtains the motion trajectories after several decoding processes.
@@ II-B2 Integrate with Transportation Context Information
The transportation context contains three types of information: intention, affordance [23], and scenario in the format of a list, in which each element is a natural language description word. For example, [“STRAIGHT”, “RIGHT-TURN”], [“SLOW-ALLOW”, “RIGHT-ALLOW”], and [“INTERSECTION”] in Fig. 1, respectively. As the number of each type of context information is fixed, we use one-hot vectors to express the context information. In addition, because the description words in the list are not exclusive, we allow several ‘1’s to appear in the one-hot vector.
More specifically, for affordance and scenario information generated by LLM, we simply encode the nature language into one-hot vectors:
@@
where $A\in\mathbb{R}^{8}$ and $S\in\mathbb{R}^{4}$ (see Fig. 1). As the number of ground truth intentions is only one, the number of description words in the intention information list is not strictly one. LLM gives possible intentions and the order of different words stands for different possibilities. So, when we encode intention information into a one-hot vector, we set different weights for different intentions in the list:
@@
where $I\in\mathbb{R}^{5}$ (see Fig. 1). So that the first word in the list has the maximal weight, and the last word holds the minimal weight.
@@
TABLE I: Effects of Different Components in Transportation Context Generation Prompt.
@@
Sugg: Prompt Design Suggestions. TCGP-Opt: Coloring the agents surrounding the red ego-agent in different colors to distinguish them easily and optimizing the representation of scene elements in the TC-Map and supplementing the corresponding scene description in text prompts. SPD-Add: Adding information about the speed of the ego-agent, the speeds of surrounding agents, and their relative positions and distances to the ego-agent in text prompts. ACC: The accuracy of intentions output by GPT4-V. When GPT4-V outputs multiple intentions, if at least one intention matches the ground truth without additional clarification, it is considered correct. 1st-$I$: Taking only the first intention from the list of intentions output by GPT4-V. (S,S-L,S-R) means consolidating the three intentions of STRAIGHT, STRAIGHT-LEFT, and STRAIGHT-RIGHT into a single STRAIGHT category.
@@ III Experiement
III-A Data distribution
@@
TABLE II: Data and Intention Accuracy of Three Ego-Agent Types.
@@
We created a confusion matrix for the results in Tab. II as shown in Fig. 4, where we have merged STRAIGHT_LEFT and STRAIGHT_RIGHT into STRAIGHT. In this matrix, the rows represent the Ground Truth intentions, while the columns represent the intentions predicted by GPT4-V. Observing the main diagonal of the confusion matrix, it’s not hard to see that STRAIGHT has the highest prediction accuracy, followed by LEFT_TURN and RIGHT_TURN predictions. The lower prediction accuracy for STATIONARY is primarily because it can be correctly identified only when the ego-agent’s speed is zero. Nevertheless, its Ground Truth is classified as STATIONARY even at very low speeds. Ego-agents intended for LEFT_U_TURN are frequently classified as LEFT_TURN, a consequence of the WOMD map’s left-turn lanes occasionally accommodating straight movements, thus increasing the likelihood of LEFT_U_TURN and LEFT_TURN being predicted as STRAIGHT. Similarly, RIGHT_U_TURN ego-agents are often mistaken as STRAIGHT, and unexpectedly, they are frequently classified as STATIONARY. This anomaly primarily arises because intentions like RIGHT_U_TURN are most common among pedestrians, whose speed and walking direction can change abruptly, posing significant challenges to accurately predicting their intentions. We also present the results of MTR, even though the result of GPT-4V on Intention is lower than MTR, it can also provide information for motion prediction and improve the accuracy of MP as shown in Tab. III.
III-B Ablation Study of TCGP
We perform ablation studies on TCGB to understand the effectiveness of each component. As shown in Tab. I,
@@ III-B Ablation Study of TCGP
The data for the ablation experiment were randomly sourced from the WOMD testing set, comprising 107 TC-Maps samples. We have the following observations: 1) TCGP achieves optimal performance when fully equipped with all its components, with every module playing a role in predicting intentions. 2) TC-Map Crop enables GPT4-V to allocate more attention to the areas that should be focused on when predicting the future intentions of the ego-agent, significantly enhancing the accuracy of intention prediction. 3) R-LT is crucial because it provides the type of lane in which the ego-agent is located, and the type of lane often determines the future direction of the vehicle.
III-C LLM Augmented Motion Prediction
Dataset and metrics. We evaluate our methods and MTR on the large-scale WOMD [25]. For the task on WOMD, 1 second of history data will be provided, and 6 motion trajectories of the agents for 8 seconds in the future are needed. We utilize the official evaluation tool to calculate the evaluation metric, in which the mAP and miss rate (MR) are the most important ones in the official leaderboard  [^2] https://waymo.com/open/challenges/2023/motion-prediction/.
Data generation. There are roughly 2 million agents in WOMD and our prompt (text prompt and image prompt) costs about \$0.1 for each agent. So generating the transportation context information for all agents in WOMD costs nearly \$200,000. As it is costly to use GPT4-V for each agent in WOMD, we first generate the transportation context information from a minimal part of WOMD and then generate similar transportation context information for the remaining part via the nearest neighbor algorithm (see Alg.1). We first split the overall dataset into two different parts: $T_{1}$ and $T_{2}$, LLM will generate transportation context information for $T_{2}$. Secondly, we encode the scenario information for each agent in $T_{1}$ and $T_{2}$ and obtain their feature vectors. Finally, for each agent $A_{i}$ in $T_{1}$, the similar transportation context is the same as the agent $A_{j}$ in $T_{2}$ that has the closest distance with $A_{i}$. More specifically, we use Euclidean distance in the nearest neighbor finding algorithm.

### v3

@@ Abstract
Motion prediction is among the most fundamental tasks in autonomous driving. Traditional methods of motion forecasting primarily encode vector information of maps and historical trajectory data of traffic participants, lacking a comprehensive understanding of overall traffic semantics, which in turn affects the performance of prediction tasks. In this paper, we utilized Large Language Models (LLMs) to enhance the global traffic context understanding for motion prediction tasks. We first conducted systematic prompt engineering, visualizing complex traffic environments and historical trajectory information of traffic participants into image prompts—Transportation Context Map (TC-Map), accompanied by corresponding text prompts. Through this approach, we obtained rich traffic context information from the LLM. By integrating this information into the motion prediction model, we demonstrate that such context can enhance the accuracy of motion predictions. Furthermore, considering the cost associated with LLMs, we propose a cost-effective deployment strategy: enhancing the accuracy of motion prediction tasks at scale with 0.7% LLM-augmented datasets. Our research offers valuable insights into enhancing the understanding of traffic scenes of LLMs and the motion prediction performance of autonomous driving. The source code is available at https://github.com/AIR-DISCOVER/LLM-Augmented-MTR and https://aistudio.baidu.com/projectdetail/7809548.
Figure 1: Context-aware Motion Prediction Based on LLMs. We first visualize the structured information of one scenario in the motion prediction dataset, GPT4-V then understand the scenario via a visualized image and well-designed prompt. Finally, GPT4-V outputs transportation context information. This information will be used to augment traditional motion prediction algorithms.
@@ I INTRODUCTION
Motion prediction is one of the most important tasks in the field of autonomous driving [1, 2, 3], which predicts the motion statuses of nearby agents by jointly considering nearby agents and road maps. This information will assist the decision module in making a more robust and safer driving decision. Therefore, this field already features a plethora of datasets and public competitions. For instance, since 2021, Waymo has been organizing competitions in motion prediction [^1] https://waymo.com/open/challenges/2021/motion-prediction/, attracting models and algorithms that have won championships, such as MTR [1], MTR++ [2], and MGTR [4]. There are three classes of motion prediction methods: goal-based methods [5, 6], direct-regression methods [7, 8], and methods that take the best of both[1, 2]. Although methods that take the best of both can balance results and computing resources greatly, they still encode information of agents and maps first and then decode to obtain final results. The methods above all lack a comprehensive understanding of overall traffic semantics.
@@
Large Language Model (LLM) is popular in the field of autonomous driving, as not only the Language Model but also the World Model with common sense [9, 10, 11]. After the release of OpenAI’s GPT series of large language models, many works that use GPT to assist driving have been emerging [12, 13, 14, 15, 16]. These efforts leverage the inferential capabilities of large language models or through experimentation, aiding autonomous driving algorithms in better perception [17] and planning [12, 14, 15], and even facilitating end-to-end decision-making [18, 13]. However, teaching LLM to understand the interactive context among various transportation participants in diverse transportation scenarios is still a challenging task. Although many have applied vision-language models to traffic scenarios, their performance heavily relies on datasets annotated from a first-person perspective [19, 20, 21]. For abstract, information-overwhelming, and bird’s-eye view (BEV) traffic scenarios, there is a lack of relevant datasets and understanding methodologies.
@@
In this paper, we propose a new method that enhances the traffic context understanding of motion prediction models using LLMs to make more accurate predictions. To enable an unfine-tuned LLM with vision (e.g. GPT4-V) to understand traffic context and output the necessary context information for motion prediction, we visualize the vector map data and historical data of traffic participants as a Transportation Context Map (TC-Map) to serve as an image prompt and propose a corresponding text prompt design. After successfully obtaining traffic context information generated by the LLM, we integrated this information into a classical motion prediction algorithm—MTR [1]. The results indicate that this context information effectively improves the accuracy of motion prediction. Additionally, considering the cost of LLMs, we also propose a cost-effective deployment strategy: by utilizing 0.7% LLM-enhanced datasets, we can empower motion prediction performance at scale.
@@
  • We systematically designed and conducted prompt engineering to enable an unfine-tuned GPT4-V to comprehend complex traffic scenarios involving multiple traffic participants and to output context information such as intention, affordance, and scenario;
@@
Figure 2: Details of Transportation Context Generation Prompt.
@@ II Method
II-A Get Transportation Context Information from GPT4-V
Translating images directly into motion poses is significantly challenging [22]. We leverage vision-enhanced LLMs to derive transportation context, identifying the agent of interest in the TC-Map as the ego-agent, categorized into vehicles (V), pedestrians (P), and cyclists (C). Illustrated in Fig. 2, the TC-Map designed based on the Waymo Open Motion Dataset (WOMD) [23] serves as the rasterized image prompt, and together with the text prompt forms the Transportation Context Generation Prompt (TCGP). The caption of the text prompt is obtained by filling in the template with code. Utilizing GPT4-V with this combined input facilitates the extraction of comprehensive context including situation understanding, reasoning, intentions of ego-agent, affordances [24], and scenario types. Additionally, six prompt design suggestions have been summarized for input prompts as shown in Fig. 3.
@@ II-A1 TC-Map Design Suggestions
Informed by prompt design suggestion 1 and 2. Suggestion 1 minimizes information overload. Drawing inspiration from the MTR Map Collection module [1], we introduced TC-Map Crop to focus the ego-agent’s attention solely on relevant surrounding scenes. Initial GPT4-V tests revealed that dense data, like 11 frames of historical trajectory in JSON format, leading to unclear intentions. Thus, we customized TC-Map dimensions: 120m x 120m for vehicles, 80m x 80m for pedestrians, and 60m x 60m for cyclists, respectively, to filter out extraneous influences on intention discernment. Furthermore, to address the varying orientations of the ego-agent across different scenarios, suggestion 2 proposes normalizing the heading of the ego-agent to “North” with a common map guide icon to reduce confusion in direction identification.
@@ II-A2 Text Prompt Design Suggestions
Addressing design suggestions 3 and 4, we introduce a structured approach to example creation. Suggestion 3 involves categorizing examples into positive and negative instances. Positive examples ensure a diverse and random generation of intentions. Negative examples are bad cases selected from GPT4-V’s responses during the quantitative experiment of prompts, focusing on scenarios with high misjudgment rates and uncommon, easily overlooked corner cases. Additional error analysis is provided and placed closer to the end of the negative example. Suggestion 4 consists of two sentences, the first sentence contains the ego-agent’s driving lane type in TC-Map determined by a rule-based algorithm, and the second sentence provides explanations for the added lane type. They aids GPT4-V in ruling out implausible intentions by assuming compliance with traffic regulations.
@@ II-A3 Detailed Information Supplement
Building on prompt design suggestion 5, we visualize more scenario elements using WOMD’s comprehensive scenario details, including road edges and crosswalks. For agents, we optimize the visualization of different scenario elements in the TC-Map by adding indicative head arrows and visual labels. Moreover, this suggestion includes integrating dynamic elements such as the ego-agent’s speed and the relative speeds, positions, and distances of nearby agents, enriching the context for more accurate motion prediction.
@@ II-A4 Multi-task Outputs Requirement
Matching to prompt design suggestion 6, we delineated seven intention categories for the ego-agent inspired by the official evaluation tool of Waymo [23]. STRAIGHT-LEFT and STRAIGHT-RIGHT anticipate potential lane changes, indicating the complexity of intentions. Predicting these intentions often requires GPT4-V to utilize rasterized images, the ego-agent’s position, heading, and speed, the state of surrounding vehicles, and the direction of nearby lane lines predicting these two intentions is highly challenging. Additionally, our approach includes the creation of five affordance types and four scenario types based on an understanding of comprehensive scenario analyses. These three types of information collectively provide rich, human-like transportation context information to enhance the accuracy of the motion prediction task.
@@ II-B1 Motion Transformer
We selected the Motion Transformer (MTR) [1], the state-of-the-art model in the WOMD Challenge 2022, as our foundational model. Leveraging the Transformer architecture [25], MTR encodes comprehensive scenario data—including agent movements and map features—before decoding this information to predict motion trajectories. This process underpins our model’s ability to accurately anticipate future positions.
@@ II-B2 Integrate with Transportation Context Information
To enrich our motion prediction model with transportation context, we categorized context information into three types: intention, affordance, and scenario, each represented as natural language descriptors (e.g. [Straight, Right-Turn], [Slow-Allow, Right-Allow], and [Intersection] in Fig. 1). These descriptors are encoded as one-hot vectors, allowing for multiple active elements to reflect the non-exclusive nature of these descriptors. Notably, affordances and scenarios are directly encoded, while intentions receive weighted encoding to prioritize them.
More specifically, for affordance and scenario information generated by LLM, we simply encode the description words into one-hot vectors:
@@
where $A\in\mathbb{R}^{8}$ and $S\in\mathbb{R}^{4}$ (see Fig. 1). LLM gives all possible intentions and the order of different description words stands for different possibilities. So, when we encode intention information into a one-hot vector, we set different weights for different intentions in the list:
@@
where $I\in\mathbb{R}^{5}$ (see Fig. 1), and each intention’s weight inversely correlates with its position in the list, emphasizing more probable intentions. So that the first word in the list has the maximal weight, and the last word holds the minimal weight.
@@
TABLE I: Effects of Different Components in TCGP
@@
Sugg: Prompt design suggestion. TCGP-Opt: Coloring agents surrounding the ego-agent in different colors to distinguish them easily, optimizing the visualization of scenario elements in the TC-Map and supplementing the corresponding scenario description in the text prompt. SPD-Add: Adding information about the speed of the ego-agent, the speeds of surrounding agents, and their relative positions and distances to the ego-agent in the text prompt. ACC: The accuracy of intentions output by GPT4-V. When GPT4-V outputs multiple intentions, if at least one intention matches the ground truth without additional clarification, it is considered correct. (1st-$I$): Taking only the first intention from the list of intentions output by GPT4-V. (S,S-L,S-R) means consolidating the three intentions of STRAIGHT, STRAIGHT-LEFT, and STRAIGHT-RIGHT into a single STRAIGHT category.
@@ III Experiement
III-A TCGP Generated Transportation Context Information
@@
TABLE II: Data and Intention Accuracy of Three Ego-Agent Types
@@
Data distribution. The data set used to generate transportation context information was randomly selected from the WOMD validation set, consisting of approximately 15,000 TC-Map samples distributed according to a vehicle (V) to pedestrian (P) to cyclist (C) ratio of 9:3:3. We create a confusion matrix for the results of intention in Tab. II as shown in Fig. 4. For simplification, the categories STRAIGHT-LEFT and STRAIGHT-RIGHT were amalgamated into a single STRAIGHT category. The matrix’s rows denote the actual intentions, while its columns represent the intentions as predicted by GPT4-V. Observing the main diagonal of the confusion matrix, it’s not hard to see that STRAIGHT has the highest prediction accuracy, followed by LEFT-TURN and RIGHT-TURN predictions. The comparatively lower accuracy observed for STATIONARY intentions stems from its correct identification only at a zero speed of the ego-agent, despite being labeled as STATIONARY at minimal speeds. Ego-agents intended for LEFT-U-TURN are frequently classified as LEFT-TURN, a consequence of the WOMD map’s left-turn lanes occasionally accommodating straight movements, thus increasing the likelihood of LEFT-U-TURN and LEFT-TURN being predicted as STRAIGHT. Similarly, RIGHT-U-TURN ego-agents are often mistaken as STRAIGHT, and unexpectedly, they are frequently classified as STATIONARY. This anomaly primarily arises because intentions like RIGHT-U-TURN are most common among pedestrians, whose speed and walking direction can change abruptly, reflecting the challenge of precisely capturing the dynamic intentions of pedestrians. Despite GPT4-V’s intention prediction accuracy being lower than that of the MTR model, the integration of GPT4-V’s output still contributes valuable context for motion prediction, enhancing overall prediction accuracy, as evidenced in Tab. III.
@@
Ablation Study. We perform ablation studies on TCGP to understand the effectiveness of each component of our prompt in Tab. I. The data comprising 107 TC-Maps samples for the ablation experiment were randomly sourced from the validation set of WOMD. We have the following observations: 1) TCGP achieves optimal performance when fully equipped with all its components, with every module playing a role in predicting intentions. 2) TC-Map Crop enables GPT4-V to allocate more attention to the areas that should be focused on when predicting the future intentions of the ego-agent, significantly enhancing the accuracy of intention prediction. 3) R-LT is crucial because it provides the type of lane in which the ego-agent is located, and the type of lane often determines the future direction of the ego-agent. These observations underscore the multifaceted contributions of TCGP’s components to the overall effectiveness of intention prediction, highlighting the importance of a holistic approach in optimizing predictive performance.
III-B LLM Augmented Motion Prediction
Dataset and metrics. We evaluated our LLM-augmented motion prediction methods using the extensive WOMD dataset. Task requirements include predicting 6 future motion trajectories over 8 seconds, based on 1 second of historical data. The official WOMD evaluation tool calculates key metrics, notably the mean Average Precision (mAP), which serve as crucial benchmarks in the official leaderboard .
Data generation. To manage the high cost of generating transportation context information for WOMD’s 2 million agents (approximately \$0.1 per agent, totaling nearly \$200,000), we implemented a cost-effective strategy. Initially, LLM-generated transportation context for a small subset of WOMD was expanded to the remainder via a nearest neighbor algorithm, utilizing Euclidean distance for similarity assessment (see Alg. 1). This approach efficiently scales LLM augmentation while maintaining data integrity.